2017/3/5(Mon.)

 * Remote participants by eZuce: Izumi

 * Report of work (*:responsible person)

  * Alignment: Enomoto, Miyo, Yokozawa
   * We found beam at the IYA but it is not clear PR or BS is misaligned.

  * IOO:
   * IMMT2 cannot be seen by camera. We will work on camera this week.

  * CRY: mirror temperature 71K.

 * Work plan (*:responsible person)

  * Alignment: Enomoto
   * We will find beam at the ETMY or ETMX by dithering PR3.

  * BS: Mark
   * We lift up the optical table and change items to vacuum compatible one. It does not change BS alignment an oplev.


 * Laser hazard: AM: PSL, PM: Upstream SR2-3 line, and Y-end.

 * Weekly
  * Week of 2/26
   * ETMY tuninng.
   * Beam alignment to REFL port.

  * week of 3/5
   * IYC door will be closed.
   * Exchanging items and cleaning inside the PR2 chamber

  * week of 3/12
   * TMS installation at ETMY.
   * Connecting ducts at the corner area.

  * week of 3/19
   * Vacuum test at at the corner area.
   * TMS alignment using green light at ETMY.


 * Issues